OpenStudy (anonymous):

Find h′(2) if h(x)=ln(x+sqrt(x^2−1√).

OpenStudy (anonymous):

i got 0.3452 this is not right

OpenStudy (anonymous):

\[h'=\frac{ 1- \frac{ 2x }{ x^2-1 } }{ x+\sqrt{x^2-1} }\]

OpenStudy (anonymous):

i don't know how you got the numerator

OpenStudy (anonymous):

waıt this\[1-\frac{ x }{ 2\sqrt{x^2-1} }\]

OpenStudy (anonymous):

yo get thıs when you derıve \[x+\sqrt{x^2-1}\]

OpenStudy (anonymous):

\[h'(x)=\frac{ 1-\frac{ x }{ 2\sqrt{x^2-1}} }{ x+\sqrt{x^2-1} }\]

OpenStudy (anonymous):

since\[y=\ln x\] \[y'=\frac{ x' }{ x }\]

OpenStudy (anonymous):

in my numerator i got \[1+\frac{ 1 }{ 2\sqrt{x^2-1} }\]

OpenStudy (anonymous):

how did you get a x where i have a 1?

OpenStudy (anonymous):

you are correct but there ıs no 2,but x on top

OpenStudy (anonymous):

where does the x come from?

OpenStudy (sirm3d):

chain rule for (x^2 - 1)

OpenStudy (anonymous):

\[\sqrt{x ^{2}-1})'= 1/2(2x)(x ^{2}-1 )^{-1/2} =\frac{ x }{ \sqrt{x ^{2}-1} }\]

OpenStudy (sirm3d):

chain rule doesn't stop on the square root; it extends to the expression inside the square root.

OpenStudy (anonymous):

ok that makes sense thanks for explaining it

OpenStudy (sirm3d):

going on, \[h\prime (2)\] equals what value?

OpenStudy (anonymous):

0.57735

OpenStudy (sirm3d):

right. or \[h\prime (2) = \frac{ \sqrt{3} }{ 3 }\]